OpenStudy (anonymous):

(3t^2+3t)/(6t^2+9t) simplify by removing factors of 1

OpenStudy (anonymous):

Factor both the numerator and the denominator. 3t is the common factor in both.\[\frac{3 t^2+3 t}{6 t^2+9 t}=\frac{(3 t (t+1))}{(3 t (2 t+3))}=\frac{t+1}{2 t+3} \]The result above is in the most simple form that I can see.